What is the solutions of x and y for the equations x+y=1 and 2x-y=-2?
Answer by ewatrrr(24785)

 
Hi,
What is the solutions of x and y for the equations
x+y=1 y+=+1-x
2x-y=-2 ||Adding to eliminating y
3x = -1
x = -1/3 and y = 4/3